# EmojiVoice: Towards long-term controllable expressivity in robot speech

### v1

@@ III-A Text-to-speech (TTS) model
We have adapted our text-to-speech model, Matcha-TTS, for the social robotics community. For the first time, we present a multi-speaker/voice version of Matcha-TTS [11] [^5] https://github.com/shivammehta25/Matcha-TTS/tree/main that allows for fast switching between emoji styles. Behind the scenes, Matcha-TTS is a neural TTS trained using optimal-transport conditional flow matching (OT-CFM). We have chosen this model for our toolbox as it is probabilistic rather than auto-regressive. Using a probabilistic model allows us to avoid hallucinations, i.e. ensuring the output audio will always match the input text, as well as issues with speaker identity control. Moreover, the model is ideal for robot’s applications as it has a compact memory footprint of only 20.9M parameters, in comparison to 830M for VoiceCraft and 41.2M for Fastspeech 2, and has a real-time-factor (RTF) of 0.3 (an RTF $<1$, taking at most 1 second to generate 1 second of audio, is considered real-time [27]) on a Nvidia Jetson AGX Orin 64GB, a GPU with similar power and memory to those applied in production robotics. Multi-speaker Matcha-TTS takes a label, in the case of the EmojiVoice toolbox an emoji number, along with the input text. This is concatenated both to the text encoder input and the flow-prediction network (decoder) input in order to contain multiple voices in a single, small (78MB, in our case) checkpoint. The architecture can be seen in Fig. 2. Having multiple voices in a single checkpoint both reduces space and allows for efficient switching between voice styles without needing to reload the model.

### v2

@@ III-A Text-to-speech (TTS) model
We have adapted our text-to-speech model, Matcha-TTS, for the social robotics community. For the first time, we present a multi-speaker/voice version of Matcha-TTS [11] [^5] https://github.com/shivammehta25/Matcha-TTS/tree/main that allows for fast switching between emoji styles. Behind the scenes, Matcha-TTS is a neural TTS trained using optimal-transport conditional flow matching (OT-CFM). We have chosen this model for our toolbox as it is probabilistic rather than auto-regressive. Using a probabilistic model allows us to avoid hallucinations, i.e. ensuring the output audio will always match the input text, as well as issues with speaker identity control. Moreover, the model is ideal for robot’s applications as it has a compact memory footprint of only 20.9M parameters, in comparison to 830M for VoiceCraft and 41.2M for Fastspeech 2, and has a real-time-factor (RTF) of 0.3 (an RTF <1<1< 1, taking at most 1 second to generate 1 second of audio, is considered real-time [27]) on a Nvidia Jetson AGX Orin 64GB, a GPU with similar power and memory to those applied in production robotics. Multi-speaker Matcha-TTS takes a label, in the case of the EmojiVoice toolbox an emoji number, along with the input text. This is concatenated both to the text encoder input and the flow-prediction network (decoder) input in order to contain multiple voices in a single, small (78MB, in our case) checkpoint. The architecture can be seen in Fig. 2. Having multiple voices in a single checkpoint both reduces space and allows for efficient switching between voice styles without needing to reload the model.